# Corrected-Inverse-Gaussian First-Hitting-Time Modeling for Molecular Communication Under Time-Varying Drift

### v1

@@ Abstract
This paper develops a tractable analytical channel model for first-hitting-time molecular communication systems under time-varying drift. While existing studies of nonstationary transport rely primarily on numerical solutions of advection–diffusion equations or parametric impulse-response fitting, they do not provide a closed-form description of trajectory-level arrival dynamics at absorbing boundaries. By adopting a change-of-measure formulation, we reveal a structural decomposition of the first-hitting-time density into a cumulative-drift displacement term and a stochastic boundary-flux modulation factor. This leads to an explicit analytical expression for the Corrected-Inverse-Gaussian (C-IG) density, extending the classical IG model to strongly nonstationary drift conditions while preserving constant-complexity evaluation. High-precision Monte Carlo simulations under both smooth pulsatile and abrupt switching drift profiles confirm that the proposed model accurately captures complex transport phenomena, including phase modulation, multi-pulse dispersion, and transient backflow. The resulting framework provides a physics-informed, computationally efficient channel model suitable for system-level analysis and receiver design in dynamic biological and molecular communication environments.
@@ I Introduction
Molecular communication (MC) channels are governed by the stochastic transport of signaling molecules through diffusion and advection [11, 9]. The dominant modeling paradigm in the literature characterizes such channels via concentration-based impulse responses [16] derived from advection–diffusion partial differential equations [10]. Under steady and uniform drift conditions, these formulations admit compact analytical expressions, and for absorbing receiver settings, the first-hitting-time (FHT) distribution provides a natural description of molecule arrival dynamics. In particular, when drift is constant, the FHT density follows the classical Inverse-Gaussian (IG) distribution [18], which has been widely adopted as a tractable baseline analytical channel model.
In realistic biological and engineered environments, however, transport conditions are rarely stationary. Pulsatile cardiovascular flows [4, 20] and time-varying electrophoretic transport in microfluidic platforms [5, 6] induce explicitly time-dependent drift velocities. Physically, these environments are commonly modeled using oscillatory velocity profiles superimposed on a nonzero mean drift, reflecting periodically driven pressure gradients in vascular systems [19, 12]. Such nonstationary drift reshapes molecule arrival statistics, leading to phase modulation, multi-peak behavior, and transient backflow effects that cannot be captured by stationary channel models.
Despite extensive investigation of time-varying transport, most existing MC studies remain concentration-centric. While recent analytical advances have successfully modeled time-variant MC channels induced by the random Brownian mobility of transceivers [3], these approaches fundamentally assume a static fluid medium governed by isotropic diffusion. Characterizing the nonstationarity induced by explicitly time-varying fluid drift remains a distinctly different and open analytical challenge. Unlike transceiver mobility, which can be resolved by statistically averaging a static channel impulse response over random distances, time-varying advection continuously alters the underlying stochastic trajectory. This triggers directional phenomena, such as transient backflow, that purely diffusion-based mobility models cannot capture. Consequently, time variability in drift is typically handled via numerical solutions of advection–diffusion partial differential equations (PDEs) or simulation-calibrated channel impulse responses (CIRs), where system parameters are adjusted dynamically to fit observed data [13]. While such approaches accurately capture macroscopic concentration evolution, they do not yield closed-form analytical models for trajectory-level arrival statistics at absorbing boundaries [1]. From a stochastic-process viewpoint, the exact FHT density under time-varying drift can in principle be characterized through Volterra-type integral equations arising from first-passage theory [7, 17]. However, these formulations generally lack closed-form solutions and require recursive numerical evaluation, making them unsuitable for real-time channel modeling and signal processing applications.
@@
To address this challenge, this paper develops a physics-informed analytical channel model for FHT behavior under time-varying drift. By reformulating stochastic transport through a change-of-measure perspective, we uncover a structural decomposition of the FHT density into a cumulative-drift displacement term and a stochastic boundary-flux modulation factor. This leads to an explicit analytical Corrected-Inverse-Gaussian (C-IG) density formula that extends the classical IG model to strongly nonstationary drift conditions while preserving constant-complexity evaluation.
The main contributions of this work are summarized as follows.
@@
  • Explicit Analytical Corrected-IG Density: We derive an explicit analytical C-IG density that generalizes the classical IG model (see [18]) to strongly nonstationary regimes. The resulting formulation captures complex phenomena—including phase modulation, multi-pulse dispersion, and transient backflow—while maintaining $\mathcal{O}(1)$ computational complexity, making it suitable for real-time applications.
  • Robustness Across Diverse Drift Profiles: Through high-precision Monte Carlo validation, we demonstrate the robustness of the C-IG model across qualitatively different scenarios, ranging from smooth periodic to abrupt switching drift. This confirms the model’s waveform-agnostic nature and establishes it as a computationally efficient alternative to PDE-based simulations for system-level analysis in molecular communication.
@@ II-A Girsanov Three-Factor Decomposition
Under the target measure $\mathbb{Q}$, the dynamics incorporate a deterministic time-varying drift, $dX_{t}=\mu(t)dt+\sigma dW_{t}.$ We assume that $\mu(t)$ is deterministic and square-integrable on finite intervals. The stopping time to an absorbing boundary $\ell>x_{0}$ is defined as $T:=\inf\{t>0:X_{t}=\ell\}.$
@@ II-B Extraction of the Macroscopic Exponential Core
Figure 1: Schematic illustration of the 1D molecular communication system under pulsatile flow. Information molecules are released by the Tx nanomachine at $x_{0}$ and propagate through a fluid medium characterized by a constant diffusion coefficient $\sigma^{2}$ and a time-varying drift velocity $\mu(t)$. The channel impulse response is determined by the First-Hitting Time (FHT) $T$, the random instant when the stochastic trajectory $X_{t}$ first contacts the absorbing boundary at $x=\ell$.
To obtain the FHT density under time-varying drift, we evaluate the conditional expectation of the Radon–Nikodym derivative given $T=t$. The principal analytical difficulty arises from the stochastic coupling term (see Eq. (1)), which involves an infinite-dimensional path integral.
To overcome this, we employ a most-probable-path approximation [8]. This corresponds to a saddle-point evaluation of the path integral in the small-diffusion regime. Conditioned on $T=t$, the dominant trajectory is approximated by the linear interpolation
@@
Substituting this path into the stochastic coupling term yields
$\frac{1}{\sigma^{2}}\int_{0}^{t}\mu^{\prime}(s)\bar{X}_{s}ds=\frac{1}{\sigma^{2}}\left[\frac{\lambda M(t)}{t}-(\mu(t)\ell-\mu(0)x_{0})\right],$ (5)
where $\lambda=\ell-x_{0}$.
@@
While this exponential core accurately captures macroscopic phase shifts of the arrival pulse, determining the precise instantaneous amplitude requires a more refined treatment of boundary flux dynamics. This amplitude correction is addressed in the following section.
@@
Figure 2: Methodological flowchart of the proposed framework. The approach leverages a change-of-measure formulation to decompose nonstationary transport into deterministic cumulative displacement and stochastic flux modulation, synthesized via the EPF correction into the final analytical C-IG model.
@@ III Instantaneous Flux Modulation and Expected Positive Flux
While the exponential core in (6) accurately captures the macroscopic timing shifts induced by the cumulative displacement $M(t)$, completely characterizing the FHT density requires determining the instantaneous amplitude prefactor. A naive tangent approximation would scale this amplitude by a deterministic instantaneous flux. However, in highly pulsatile regimes, strong backflow ($\mu(t)<0$) causes deterministic models to incorrectly predict zero or negative arrival rates, leading to artificial truncation. Since thermal diffusion physically guarantees a nonzero boundary-crossing probability even against macroscopic backflow, the instantaneous boundary flux must be modeled as an inherently stochastic quantity. This motivates our diffusion-consistent Expected Positive Flux (EPF) formulation.
@@ III-B Expected Positive Flux Formulation
To resolve the non-physical truncation during backflow phases, we model the instantaneous boundary flux as a normally distributed random variable. The macroscopic flux $F_{\text{mean}}(t)$ acts as the mean, while the local diffusion scale $S(t)=\sqrt{\sigma^{2}t}$ acts as the standard deviation.
The effective arrival rate is governed by the expectation of the strictly positive components of this flux distribution. By defining the standardized flux state $Z(t)=F_{\text{mean}}(t)/S(t)$, the smooth Expected Positive Flux $F_{\text{smooth}}(t)$ is derived as
@@
The EPF operator in (8) acts as a rigorous, physics-informed “soft-plus” function. During forward flow stages ($Z\gg 0$), $F_{\text{smooth}}\approx F_{\text{mean}}$, perfectly preserving the heightened peak arrivals. Conversely, during severe backflow ($Z<0$), the EPF yields a smooth, strictly positive tail driven purely by the diffusion term $S(t)\phi(Z)$, accurately preventing the density from collapsing to zero.
@@ III-C The Corrected-Inverse-Gaussian Channel Model
By combining the exponential core derived in (6) with EPF formulation in (8), we obtain an explicit analytical characterization of the FHT density under time-varying drift.
Result (Corrected-Inverse-Gaussian Density). Under deterministic time-varying drift $\mu(t)$ and diffusion coefficient $\sigma^{2}$, the FHT density at an absorbing boundary $\ell$ is approximated by the C-IG formula,
@@ IV-A Validation Under Strongly Pulsatile Drift
Fig. 3 compares the FHT distributions. The classical IG model, which assumes constant drift $v_{0}$, fails to capture the oscillatory modulation of arrival density and substantially overestimates peak amplitudes. In contrast, the proposed C-IG density closely matches the Monte Carlo results across all time regimes. The cumulative displacement $M(t)$ accurately predicts phase-shifted arrival peaks, while the EPF mechanism preserves nonzero arrival probability during transient backflow intervals.
These results demonstrate that the C-IG framework extends the analytical tractability of the IG structure to strongly time-varying drift conditions, while retaining constant-complexity evaluation.
Figure 3: First-hitting-time distributions under sinusoidal pulsatile drift. The proposed C-IG density captures phase shifts and amplitude modulation, whereas the classical IG model fails to represent nonstationary transport effects.
@@ IV-B Robustness Under Abrupt Drift Switching
This profile introduces a discontinuity in the drift derivative and produces large total variation in the driving field. Under such abrupt switching, the classical IG model exhibits severe peak misalignment and amplitude distortion.
In contrast, the C-IG density remains stable. The cumulative displacement $M(t)$ correctly predicts the macroscopic phase transition at the switching time, and the EPF prefactor prevents artificial truncation during pullback intervals. As shown in Fig. 4, the analytical density reproduces both the sharp peak formation and the post-switch depletion observed in Monte Carlo simulations.
Figure 4: First-hitting-time distributions under single-step drift switching with $T_{sw}=1.5$. The proposed C-IG density remains accurate under abrupt drift variation, while the classical IG model fails to align with the observed phase transition.
@@ IV-C Model Limitations and Practical Implications
While the C-IG model exhibits strong agreement with Monte Carlo simulations, minor deviations appear in later arrival peaks under highly nonstationary drift. These discrepancies arise from irreversible absorption effects—higher-order memory phenomena where early-arriving particles are permanently removed, gradually skewing the surviving population in a manner not captured by the C-IG density. Nevertheless, these deviations remain secondary compared to the substantial mismatch of the stationary IG baseline (see Fig. 3). By accurately capturing phase modulation and amplitude restructuring while preserving constant-complexity evaluation, the C-IG framework remains well-suited for system-level analysis, receiver optimization, and real-time channel estimation in nonstationary MC systems.
@@ V Conclusion
This paper established a tractable analytical framework for modeling FHT dynamics under nonstationary drift. By leveraging a change-of-measure perspective, we derived the C-IG density (see Eq. (9))—an explicit analytical expression that generalizes the classical IG model to dynamic environments. Particle-level simulations confirmed that the C-IG model accurately captures complex transport phenomena, including phase modulation, multi-pulse dispersion, and transient backflow, while maintaining constant evaluation complexity. This framework bridges the gap between physics-faithful transport modeling and tractable channel analysis, providing a robust foundation for the design and optimization of nonstationary biological and MC systems.
Acknowledgement

### v2

# The Corrected Inverse-Gaussian: A Tractable First-Hitting-Time Channel Model for Nonstationary Molecular Communication

@@ Abstract
This paper develops a tractable analytical channel model for first-hitting-time molecular communication (MC) systems under time-varying drift. While existing studies of nonstationary transport rely primarily on numerical solutions of advection–diffusion equations or parametric impulse-response fitting, they do not provide an explicit analytical description of trajectory-level arrival dynamics at absorbing boundaries. By adopting a change-of-measure formulation, we reveal a structural decomposition of the first-hitting-time density into a cumulative-drift displacement term and a stochastic boundary-flux modulation factor. This leads to a closed-form Corrected-Inverse-Gaussian (C-IG) density that generalizes the classical IG model to nonstationary drift while preserving $\mathcal{O}(1)$ evaluation complexity. Monte Carlo simulations under both smooth pulsatile and abrupt switching drift profiles confirm that the proposed C-IG model accurately captures complex transport phenomena, including phase modulation, multi-pulse dispersion, and transient backflow—effects that traditionally complicate symbol synchronization and induce severe inter-symbol interference. The resulting framework provides a physics-informed, computationally efficient channel model suitable for system-level analysis and advanced receiver design, such as real-time maximum likelihood detection, in dynamic biological and MC environments.
@@ I Introduction
Diffusion-based molecular communication (MC) channels are governed by the stochastic transport of signaling molecules through diffusion and advection [11, 9]. The dominant modeling paradigm characterizes such channels via concentration-based impulse responses [16] derived from advection–diffusion partial differential equations (PDEs) [10]. Under steady and uniform drift conditions, these formulations admit compact analytical expressions [11]. For absorbing receiver settings, the first-hitting-time (FHT) distribution provides a natural description of molecule arrival dynamics. In particular, when drift is constant, the FHT density follows the classical Inverse-Gaussian (IG) distribution [18], which has been widely adopted as a tractable baseline analytical channel model in MC.
However, realistic transport conditions are rarely stationary. Pulsatile cardiovascular flows [4, 20] and time-varying electrophoretic transport in microfluidic platforms [5, 6] induce explicitly time-dependent drift velocities. Physically, these environments are commonly modeled using oscillatory velocity profiles superimposed on a nonzero mean drift, reflecting periodically driven pressure gradients [19, 12]. This nonstationary drift reshapes arrival statistics, causing phase modulation, multi-peak behavior, and transient backflow that stationary models cannot capture.
Despite extensive studies on time-varying transport, most MC literature remains concentration-centric. While recent advances analytically model time-variant channels induced by random transceiver mobility [3], they fundamentally assume a static fluid medium. Characterizing nonstationarity from explicitly time-varying drift remains an open analytical challenge. Unlike mobility, which is resolvable by averaging static responses over random distances, time-varying advection continuously alters the stochastic trajectory. This triggers directional phenomena, like transient backflow, that traditional models cannot capture. Consequently, drift time-variability is typically handled via numerical advection–diffusion PDEs or simulation-calibrated impulse responses [13]. While capturing macroscopic concentration, these approaches do not yield closed-form analytical models for trajectory-level arrival statistics [1]. Stochastically, the exact time-varying FHT density can be characterized through Volterra-type integral equations [7, 17]. However, lacking closed-form solutions, these recursive formulations are unsuitable for real-time channel modeling and signal processing.
@@
To address this, we develop a physics-informed analytical channel model for FHT behavior under time-varying drift. Using a change-of-measure perspective, we uncover a structural decomposition of the FHT density into a cumulative-drift displacement and a stochastic boundary-flux modulation factor. This yields a closed-form Corrected-Inverse-Gaussian (C-IG) density formula, extending the classical IG model to strongly nonstationary drift while preserving constant-complexity evaluation.
The main contributions of this work are summarized as follows:
@@
  • Closed-Form Corrected-IG Density: We derive a closed-form C-IG density that generalizes the classical IG model [18] to strongly nonstationary regimes. The resulting formulation captures complex phenomena—including phase modulation, multi-pulse dispersion, and transient backflow—while maintaining constant-time evaluation per time instance, making it suitable for real-time applications.
  • Robustness Across Diverse Drift Profiles: Through high-precision Monte Carlo validation, we demonstrate the robustness of the C-IG model across qualitatively different scenarios, ranging from smooth periodic to abrupt switching drift. This confirms the model’s waveform-agnostic nature and establishes it as a computationally efficient alternative to PDE-based simulations for system-level analysis.
@@ II-A Girsanov Three-Factor Decomposition
Under the target measure $\mathbb{Q}$, the dynamics incorporate a deterministic time-varying drift,
$dX_{t}=\mu(t)dt+\sigma dW_{t}.$ (3)
Assume that $\mu(t)$ is deterministic and square-integrable on finite intervals. The stopping time to an absorbing boundary $\ell>x_{0}$ is defined as $T:=\inf\{t>0:X_{t}=\ell\}.$
@@ II-B Extraction of the Macroscopic Exponential Core
Figure 1: Schematic illustration of the 1D molecular communication system under pulsatile flow. Information molecules are released by the Tx nanomachine at $x_{0}$ and propagate through a fluid medium characterized by a constant diffusion coefficient $\sigma^{2}$ and a time-varying drift velocity $\mu(t)$. The channel impulse response is determined by the first-hitting-time $T$, the random instant when the stochastic trajectory $X_{t}$ first contacts the absorbing boundary at $x=\ell$.
To obtain the FHT density under time-varying drift, we evaluate the conditional expectation of the Radon–Nikodym derivative given $T=t$. The principal analytical difficulty arises from the stochastic coupling term in (1), which involves an infinite-dimensional path integral. (Note that for piecewise-smooth drift profiles, the derivation applies on each smooth segment, with discontinuities treated in a distributional sense.)
To maintain tractability, we adopt a most-probable-path (MPP) approximation [8] via the linear interpolation
@@
This choice is asymptotically justified for short-time/moderate-drift regimes and, crucially, enables a critical structural cancellation that yields a constant-complexity closed-form model. Substituting this MPP path into the stochastic coupling term yields
$\begin{split}-\frac{1}{\sigma^{2}}\int_{0}^{t}\mu^{\prime}(s)\bar{X}_{s}\,ds&=-\frac{\mu(t)\ell-\mu(0)x_{0}}{\sigma^{2}}\\ &\quad+\frac{\ell-x_{0}}{\sigma^{2}t}\int_{0}^{t}\mu(s)ds.\end{split}$ (6)
@@
While this exponential core determines the macroscopic timing of arrivals, the amplitude prefactor arises from boundary-crossing dynamics, which we address next.
@@
Figure 2: Methodological flowchart of the proposed framework. The approach leverages a change-of-measure formulation to decompose nonstationary transport into deterministic cumulative displacement and stochastic flux modulation, synthesized via the EPF correction into the final closed-form C-IG model.
@@ III Instantaneous Flux Modulation and Expected Positive Flux
While the exponential core in (7) captures timing shifts through the cumulative displacement $M(t)$, the instantaneous amplitude depends on stochastic boundary-crossing dynamics. Deterministic flux approximations break down under strong backflow ($\mu(t)<0$), producing artificial truncation despite diffusion-driven arrivals. We therefore introduce the diffusion-consistent Expected Positive Flux (EPF) formulation.
@@ III-B Expected Positive Flux Formulation
To prevent non-physical truncation during backflow phases, we model the instantaneous boundary flux as a Gaussian random variable, where the macroscopic flux $F_{\text{mean}}(t)$ serves as the mean and the local diffusion scale $S(t)=\sqrt{\sigma^{2}t}$ as the standard deviation. This Gaussian approximation follows from the local central-limit scaling of diffusive fluctuations near the boundary.
The effective arrival rate is governed by the expectation of the strictly positive components of this flux distribution. By defining the standardized flux state $Z(t)=F_{\text{mean}}(t)/S(t)$, the smooth EPF $F_{\text{smooth}}(t)$ is derived as
@@
The EPF operator in (9) acts as a rigorous, physics-informed “soft-plus” function, effectively serving as a non-linear rectification mechanism for the boundary flux. During forward flow stages ($Z\gg 0$), $F_{\text{smooth}}\approx F_{\text{mean}}$, perfectly preserving the heightened peak arrivals. Conversely, during severe backflow ($Z<0$), the EPF yields a smooth, strictly positive tail driven purely by the diffusion term $S(t)\phi(Z)$, accurately preventing the density from collapsing to zero.
@@ III-C The Corrected-Inverse-Gaussian Channel Model
By combining the exponential core derived in (7) with EPF formulation in (9), we obtain a closed-form (approximated) analytical characterization of the FHT density under time-varying drift.
Main Result (Corrected-Inverse-Gaussian Density). Under deterministic time-varying drift $\mu(t)$ and diffusion coefficient $\sigma^{2}$, the FHT density at an absorbing boundary $\ell$ is approximated by the C-IG formula,
@@
Remark (Computational Complexity and Receiver Design).
A key advantage of the proposed C-IG density is its $\mathcal{O}(1)$ evaluation complexity per time instance, requiring only basic algebraic and standard normal distribution evaluations. In contrast, obtaining the exact FHT density under nonstationary drift typically requires either high-latency Monte Carlo path simulations or recursive numerical integration of Volterra-type equations [7], both of which entail significant computational overhead that scales with the desired precision. By providing a tractable $\mathcal{O}(1)$ likelihood function, the C-IG model bypasses these numerical bottlenecks, making advanced signal processing techniques, such as real-time Maximum Likelihood (ML) detection and dynamic channel equalization, practically feasible for nonstationary MC receivers.
@@ IV-A Validation Under Strongly Pulsatile Drift
Fig. 3 compares the FHT distributions. The classical IG model, which assumes constant drift $v_{0}$, fails to capture the oscillatory modulation of arrival density—a phenomenon that induces severe non-traditional inter-symbol interference (ISI)—and substantially overestimates peak amplitudes. In contrast, the proposed C-IG density closely matches the Monte Carlo results across all time regimes. The cumulative displacement $M(t)$ accurately predicts phase-shifted arrival peaks, which are critical sources of timing synchronization errors, while the EPF mechanism preserves nonzero arrival probability during transient backflow intervals.
These results demonstrate that the C-IG framework extends the analytical tractability of the IG structure to strongly time-varying drift conditions, while retaining constant-complexity evaluation, making it highly suitable for signal processing applications at the receiver.
Figure 3: First-hitting-time distributions under sinusoidal pulsatile drift. The proposed C-IG density captures phase shifts (which complicate timing synchronization) and amplitude modulation (which exacerbates ISI), whereas the classical IG model fails to represent nonstationary transport effects.
@@ IV-B Robustness Under Abrupt Drift Switching
This profile introduces a discontinuous drift derivative and large total variation in the driving field, under which the classical IG model exhibits severe peak misalignment and amplitude distortion, leading to severely suboptimal detection if directly applied to receiver design. In contrast, the C-IG density remains stable: the cumulative displacement $M(t)$ captures the macroscopic phase transition at the switching time, while the EPF prefactor prevents artificial truncation, accurately reproducing both sharp peak formation and post-switch depletion observed in Monte Carlo simulations (Fig. 4).
Figure 4: First-hitting-time distributions under single-step drift switching with $T_{sw}=1.5$. The proposed C-IG density remains accurate under abrupt drift variation, while the classical IG model fails to align with the observed phase transition, highlighting its inadequacy for reliable symbol synchronization.
@@ IV-C Model Limitations and Practical Implications
While the C-IG model exhibits strong agreement with Monte Carlo simulations, minor deviations appear in later arrival peaks under highly nonstationary drift. These discrepancies arise from irreversible absorption effects—higher-order memory phenomena where early-arriving particles are permanently removed, gradually skewing the surviving population in a manner not captured by the closed-form density. Nevertheless, these deviations remain secondary compared to the substantial mismatch of the stationary IG baseline (see Fig. 3). By accurately capturing phase modulation and amplitude restructuring while preserving constant-complexity evaluation, the C-IG framework remains well-suited for system-level analysis, receiver optimization, and real-time channel estimation in nonstationary MC systems.
@@ V Conclusion
This paper established a tractable analytical framework for modeling FHT dynamics under nonstationary drift. By leveraging a change-of-measure perspective, we derived the C-IG density (10)—a closed-form expression with $\mathcal{O}(1)$ evaluation complexity that generalizes the classical IG model to dynamic environments. Particle-level simulations confirmed that the C-IG model accurately captures complex transport phenomena, including phase modulation, multi-pulse dispersion, and transient backflow, while bypassing the latency bottlenecks of traditional numerical solvers. This framework bridges the gap between physics-faithful transport modeling and tractable channel analysis, providing a robust foundation for the signal processing design and optimization of nonstationary biological and MC receivers.
Acknowledgment